# A concise tutorial review of Reverse Osmosis and Electrodialysis

### v5

@@ 4 Theory of electrodialysis for a 1:1 salt solution
and this representation shows that this model can be used for two very different cases: first, the above-discussed geometry of an ED cell pair with two co-current flow channels, but we can also use it for an ED stack with relatively low desalination ‘per pass’, where the effluents of each channel are rerouted to two storage tanks from which the same channel is fed again. In this second type of calculation, $t^{*}$ really is a time. In that case, the area/volume ratio, $a\mathrm{{}_{d}}$, in Eq. (31) is the area of one membrane times the number of cell pairs, divided by the volume of the d-reservoir (including volumes of tubing and the channels in the ED stack). The other modification for this second type of calculation is that in Eq. (30), volume flow rates are replaced by the two reservoir volumes. Finally, we have to find expressions for current efficiency, $\lambda\mathrm{{}_{th}}$, and current density, $I$, as function of $c\mathrm{{}_{d}}$ and the applied cell pair voltage $V\mathrm{{}_{cp}}$. This is discussed in the next section.
@@ 4.1 Calculation of current and current efficiency in ED
Like Eq. (32), we have a current-voltage relationship across the d- and c-channels, given by
@@
where $\infty$ refers to a position just outside the membrane (in the ED model z-dependent), and Eq. (34) applies to the interface of the membrane with electrolyte both in the d- and c-channels. The Donnan potential at each of these interfaces is the potential just in the membrane, relative to just outside, in solution. A partitioning coefficient, $\Phi_{i}$, is also included, which we discussed in section 3.
Based on electroneutrality in the membrane, valid at each position in the membrane, we have
@@
where X without ||$\dots$|| … | can be of either sign, positive and negative. Based on Eqs. (34) and (35) we can derive for the Donnan potential that
@@
where sinh(x) is ½$\text{\textonehalf}\left(\exp\left(x\right)-\exp\left(-x\right)\right)$, and sinh-1 is the inverse function. For both ions we assume the same value of $\Phi_{i}$. As Eq. (36) shows, for a membrane with positive fixed charge, i.e., an AEM, which needs to ‘pull in the anions’, the Donnan potential is positive, i.e., from just outside to just inside the membrane, the potential goes up.
The eight voltage drops described in the above equations can all be added up (taking care of appropriate choices of $\pm$-signs), to arrive after multiplication with $V\mathrm{{}_{T}}$, at the cell pair voltage, $V\mathrm{{}_{cp}}$. And this $V\mathrm{{}_{cp}}$ is invariant with position z; see Fig. 2 in ref. [41] for an example of a voltage profile. (Thus the eight individual voltages described above, they all change with z, but their summation, which is $V\mathrm{{}_{cp}}$, does not.)
@@
which we can call the total Donnan potential across a membrane. (This excludes the membrane potential across the inner part of the membrane due to current, as given by Eq. (32).) From this point forward, we again drop index $\infty$. Now, generally, Eq. (37) is not applied, but instead, in literature what is often used is
@@
and this expression is valid for an ideal membrane, one that perfectly blocks coions. There is no dependence on membrane charge here. A prefactor in front of the ln-term can be added, which is then called the permselectivity, which describes the difference between the ideal ln-term and the measured Donnan potential. The $\pm$-symbol depends on the sign of the membrane charge and on whether ‘$\Delta$’ is defined from c-side to d-side or vice-versa.
@@
where the total Donnan potential decreases (in magnitude) when the membrane charge density X goes down, or when the difference between $c\mathrm{{}_{c}}$ and $c\mathrm{{}_{d}}$ increases. Note that this novel equation is only valid as a first correction to Eq. (38) when the salt concentrations on boh sides are less than $|X|$. For higher salt concentrations, we must use Eq. (37). Note that in Eqs. (37) and (39) it is important to use the same unit for c and for X, either both in M or both in mM, i.e., mol/m3.
@@
For the ideal model that we consider, we can now add up all eight voltages, and obtain one final equation for voltage $V\mathrm{{}_{cp}}$ as function of current density I. The result shown next assumes that c- and d-channels have the same width, $L\mathrm{{}_{ch}}$, and we already discussed the assumption that each membrane has the same charge density, $|{X}|$, and the same transport properties, thus the same $k^{*}_{\text{m}}$. We then arrive at
@@
Finally we evaluate the current efficiency $\lambda\mathrm{{}_{th}}$, at each z-position. To that end we must establish what is the flux of counterions through each membrane, which is the desired effect, leading to desalination, and the flux of coions, which is a deleterious effect, reducing desalination, because coions leak out of the c-channels to the d-channels. Based on fluxes through a single membrane, $\lambda\mathrm{{}_{th}}$ is given by
@@
where $J\mathrm{{}_{m,ions}}$ is the total ions flow through a membrane, $J\mathrm{{}_{m,ions}}=J\mathrm{{}_{m,+}}+J\mathrm{{}_{m,-}}$, and the current density is $I=F\,\left(J\mathrm{{}_{m,+}}-J\mathrm{{}_{m,-}}\right)$. The factor $\lambda\mathrm{{}_{th}}$ is only defined for desalination of a 1:1 salt. When we start with both channels fed with the same water, $\lambda\mathrm{{}_{th}}$ will be between 0 and 1, with $\lambda\mathrm{{}_{th}}=1$ for ideal operation, with only counterion transport through each membrane. In the limit of $\lambda\mathrm{{}_{th}}=1$, for each unit of charge a full salt molecule is removed from a diluate channel, while in the limit of $\lambda\mathrm{{}_{th}}=0$ current is passing the membrane, but there is no net salt removal. This last situation develops if we make an ED cell very long, or continue to recycle salt back into an ED stack long enough, see the box on p. 4.
Next, we analyze ion transport, both of counterions and coions, across a membrane. To that end we use the Nernst-Planck (NP) equation that can be solved at any position in the membrane together with local electroneutrality. For each ion, the extended NP-equation is
@@
In the derivation of Eq. (43) and (44), the electroneutrality condition in the membrane is implemented. Because of steady-state both equations can be integrated across the membrane, leading for current I to Eq. (32), and for $J\mathrm{{}_{ions,m}}$ to
$J\mathrm{{}_{ions,m}}=\pm k^{*}_{\text{m}}\left(\Delta c\mathrm{{}_{T,m}}-X\Delta\phi\mathrm{{}_{m}}\right)\,.$ (45)
where $\Delta c\mathrm{{}_{T,m}}\rangle$ is the difference between $c\mathrm{{}_{T,m}}$ at the two outsides of the membrane, which we calculate based on Eqs. (34) and (35). For a 1:1 salt, we arrive at
@@
At high enough $|X|$ (relative to $c\mathrm{{}_{c}}$ and $c\mathrm{{}_{d}}$), from Eq. (46) we can derive that the difference in $c\mathrm{{}_{T,m}}$ across the membrane can be simplified to $\Delta c\mathrm{{}_{T,m}}=2\,\Phi_{i}^{2}\,\left(c_{\text{c}}^{2}-c_{\text{d}}^{2}\right){\big{/}}\,|X|$, a result that we can implement in Eq. (47) to arrive at
@@
which is the final equation that we need so we can solve the full co-current ED model. All fluxes and concentrations in Eq. (45) depend on position z in the module, or alternatively, depend on time, or time-on-stream, $t^{*}$. Though we do not analyse that here, but the average current in a cell, as will be measured, is calculated by $\left\langle I\right\rangle=1/t^{*}\int_{0}^{t^{*}}I\text{d}t^{*}$.
@@
Eq. (48) is interesting in its own right, as it shows that $\lambda\mathrm{{}_{th}}$ will be high, near unity, at high currents and a high membrane charge, and for low salt concentration differences, but current efficiency will go down for lower currents and lower membrane charge. It also goes down when $k^{*}\text{m}$ goes up, for instance because membrane thickness is reduced. Indeed, for very thin membranes, ED does not work, because then there is a large flux of coions across the membranes and that limits the desalination process [41]. Eq. (48) illustrates that $\lambda\mathrm{{}_{th}}$ depends strongly on position in the cell, because the current density I changes with z, as well as concentrations $c\mathrm{{}_{d}}$ and $c\mathrm{{}_{c}}$.
Calculations are made with this model for co-current flow in ED for a constant $V\mathrm{{}_{cp}}$ across the cell pair (the same value of $V\mathrm{{}_{cp}}$ at each position z). Results show that while water in the d-channel is desalinated, the Donnan potential strongly increases on the diluate side (by a value which exceeds the decrease on the concentrate side), the current goes down, and current efficiency decreases, finally to reach zero, see Fig. 5. We can use this result to calculate from Eq. (48) the final salt concentration in a single-pass co-current ED cell pair when the cell is large enough, together with Eqs. (30) and (40). Calculation results are presented in Fig. 6 and show the final, i.e., minimum, $c\mathrm{{}_{d}}$ that can be reached in an ED stack as function of cell pair voltage $V\mathrm{{}_{cp}}$ for two values of water recovery, wr. Also the current density in this ‘steady-state’ limit is plotted. Calculations are based on $c\mathrm{{}_{f}}=0.5$ M, $X\!=\!\pm 4.0$ M, $k^{*}_{\text{m}}\!=\!1\text{ }\mu$m/s, $\Phi_{i}=1$, and $k\mathrm{{}_{ch}}\!=\!5\text{ }\mu$m/s.
@@
In summary, the co-current ED model is a powerful and insightful model to discuss and quantify several well-known observations in ED, such as the decrease of the current density and current efficiency with time, see Fig. 5.
@@
Figure 5: Calculations of desalination in electrodialysis with co-current flow in diluate and concentrate channels for $V\mathrm{{}_{cp}}\!=\!180$ mV, $c\mathrm{{}_{f}}\!=\!0.5$ M, $\text{{wr}}\!=\!0.5$, and $X\!=\pm\!4.0$ M. As function of time-on-stream, t*, salt concentrations in the two channels change, first rapidly, then slowly, while current I goes down as well as current efficiency, $\lambda\mathrm{{}_{th}}$. In steady-state (at large t*), there is still a flow of current across the channels and membranes, but the net salt transport is zero, thus $\lambda\mathrm{{}_{th}}\rightarrow 0$.
Figure 6: Concentrations and current density in ED based on co-current flow. For two values of wr we analyze as function of $V\mathrm{{}_{cp}}$ the steady-state diluate concentration, $c\mathrm{{}_{d}}$, which can be reached after a long enough residence time t*, as well as the resulting final current density for $\text{{wr}}\!=\!0.5$.
@@ 5.1 Introduction
For transport across the membrane, be it for membranes as thin as 100 nm as in RO, and membranes 100 to 1000 times thicker for ED, we can start with continuum theory for flow of water and ions through thin channels and capillaries, and apply this theory to transport in the pores in RO and ED membranes. These models, such as the space charge (SC) model [44, 45, 46, 47, 43], then become much more simple because profiles in concentration and potential across the cross-section of a pore in the membrane then do not need to be considered. This uniform potential (UP) model, i.e., the SC-model with averaging over the cross-section of the pore, describes ion transport and water flow through the pores of a membrane, and considers the effect of pressure on both membrane interfaces [34, 43, 48, 49]. The UP-model is thus an extension of the Teorell-Meyer-Sievers (TMS) model that describes ion transport across membranes, but TMS theory does not include water flow [50]. The UP-model can now be extended to incorporate many more effects than we could in the SC model. Besides water-membrane and water-ion friction, we can now also include ion-membrane friction and ion-ion friction, the latter describing how ions that overtake slower ones drag these along, and vice-versa slow ions will retard faster ones. This is the solution-friction (SF) model, and it describes both the flow of water and of solutes [51, 52]. In an extended version, also ion volume effects can be included, and this is called the two-fluid model (TFM) [53, 54, 55].
In SF theory, the velocity of solutes (such as ions) is determined by a balance between driving forces and frictions. Driving force acting on a solute are mathematically formulated as the negative of a gradient of an ion’s chemical potential. Frictions experienced by a solute are with the water flowing through the pores, and with the membrane matrix, which acts as an immobile background structure. For the water a similar force balance is set up but driving forces acting on the water are hydrostatic and osmotic pressures gradients [40, 38]. Thus ions and other solutes are described in a very different manner from how the water that flows past the ions through the pores is described. This is one of the key points of SF theory [56, 57, 58]. This difference in treatment is required because water is a continuum fluid and fills up all available space in the porous medium (such as a membrane), and thus forms continuous liquid pathways. Ions and other solutes are discrete, dispersed entities, and they travel across these pathways of water. These different modes of transport for water and for solutes are essential elements of SF theory

### v6

@@ 4 Theory of electrodialysis for a 1:1 salt solution
and this representation shows that this model can be used for two very different cases: first, the above-discussed geometry of an ED cell pair with two co-current flow channels, but we can also use it for an ED stack with relatively low desalination per pass, where the effluents of each channel are rerouted to two storage tanks from which the same channel is fed again. In this second type of calculation, $t^{*}$ really is a time. In that case, the area/volume ratio, $a\mathrm{{}_{d}}$, in Eq. (31) is the area of one membrane times the number of cell pairs, divided by the volume of the d-reservoir (including volumes of tubing and the channels in the ED stack). The other modification for this second type of calculation is that in Eq. (30), volume flow rates are replaced by the two reservoir volumes. Finally, we have to find expressions for current efficiency, $\lambda\mathrm{{}_{th}}$, and current density, $I$, as function of $c\mathrm{{}_{d}}$ and the applied cell pair voltage $V\mathrm{{}_{cp}}$. This is discussed in the next section.
@@ 4.1 Calculation of current and current efficiency in ED
Similar to Eq. (32), we have a current-voltage relationship across the d- and c-channels, given by
@@
where $\infty$ refers to a position just outside the membrane (in the ED model z-dependent), and Eq. (34) applies to the interface of the membrane with electrolyte both in the d- and c-channels. The Donnan potential at each of these interfaces is the potential just in the membrane, relative to just outside, in solution. A partition coefficient, $\Phi_{i}$, is also included, which we discussed in section 3. For IEMs, $\Phi_{i}$ is likely not very small, not $\ll 0.5$.
Electroneutrality in the membrane is given by
@@
where X can be positive or negative. Based on Eqs. (34) and (35) we can derive the Donnan potential
@@
where sinh(x) is ½$\text{\textonehalf}\left(\exp\left(x\right)-\exp\left(-x\right)\right)$, and sinh-1 is the inverse function. For both ions we assume here the same value of $\Phi_{i}$. As Eq. (36) shows, for a membrane with positive fixed charge, i.e., an AEM, which needs to ‘pull in the anions’, the Donnan potential is positive, i.e., from just outside to just inside the membrane, the potential goes up.
The eight voltage drops described in the above equations can all be added up (taking care of appropriate choices of $\pm$-signs), to arrive after multiplication with $V\mathrm{{}_{T}}$, at the cell pair voltage, $V\mathrm{{}_{cp}}$. And this $V\mathrm{{}_{cp}}$ is invariant with position z; see Fig. 2 in ref. [43] for an example of a voltage profile. (Thus the eight individual voltages described above, all change with z, but their summation, which is $V\mathrm{{}_{cp}}$, does not.)
@@
which we can call the total Donnan potential across a membrane. (This excludes the membrane potential across the inner part of the membrane due to current, as given by Eq. (32).) From this point forward, we drop index $\infty$ once again. Eq. (37) is not generally used, but instead, often found in literature is
@@
and this expression is valid for an ideal membrane, one that perfectly blocks coions. There is no dependence on membrane charge here. A prefactor in front of the ln-term can be added, which is then called the permselectivity, which describes the difference between the ideal ln-term and the measured Donnan potential. The $\pm$-symbol depends on the sign of the membrane charge and on whether the Donnan potential is defined from c-side to d-side, or vice-versa.
@@
where the total Donnan potential decreases (in magnitude) when the membrane charge density X goes down, or when the difference between $c\mathrm{{}_{c}}$ and $c\mathrm{{}_{d}}$ increases. This novel equation is only valid as a first correction to Eq. (38) when the salt concentrations on both sides are sufficiently lower than $|X|$. For higher salt concentrations, we must use Eq. (37). Note that in Eqs. (37) and (39) it is important to use the same unit for c and for X, either both in M or both in mM, i.e., mol/m3.
@@
For the ideal model that we consider, we can now add up all eight voltages, and obtain one final equation for voltage $V\mathrm{{}_{cp}}$ as function of current density I. The results shown next assume that c- and d-channels have the same width, $L\mathrm{{}_{ch}}$, and we already discussed the assumption that each membrane has the same charge density, $|{X}|$, and the same transport properties, thus the same $k^{*}_{\text{m}}$. We then arrive at
@@
Finally we evaluate the current efficiency $\lambda\mathrm{{}_{th}}$, at each z-position. To that end we must establish what is the flux of counterions through each membrane, which is the desired effect, leading to desalination, and the flux of coions, which reduces desalination because coions leak out of the c-channels to the d-channels. Based on fluxes through a single membrane, $\lambda\mathrm{{}_{th}}$ is given by
@@
where $J\mathrm{{}_{m,ions}}$ is the total ions flow through a membrane, $J\mathrm{{}_{m,ions}}=J\mathrm{{}_{m,+}}+J\mathrm{{}_{m,-}}$, and the current density is $I=F\,\left(J\mathrm{{}_{m,+}}-J\mathrm{{}_{m,-}}\right)$. The factor $\lambda\mathrm{{}_{th}}$ is only defined for desalination of a 1:1 salt. When we start with both channels fed with the same water, $\lambda\mathrm{{}_{th}}$ will be between 0 and 1, with $\lambda\mathrm{{}_{th}}=1$ for ideal operation, with only counterion transport through each membrane, and then for each unit of charge a full salt molecule is removed from a diluate channel. Instead, in the limit of $\lambda\mathrm{{}_{th}}=0$, current is passing the membrane, but there is no net salt removal. This last situation develops if we make an ED cell very long, or continue to recycle salt back into an ED stack for a very long time.
Next, we analyze ion transport, both of counterions and coions, across a membrane. To that end we use the Nernst-Planck (NP) equation jointly with local electroneutrality. For each ion, the extended NP-equation is
@@
In the derivation of Eq. (43) and (44), the electroneutrality condition in the membrane is implemented. Because of steady state, both equations can be integrated across the membrane, leading for current I to Eq. (32), and for $J\mathrm{{}_{ions,m}}$ to
$J\mathrm{{}_{ions,m}}=\pm k^{*}_{\text{m}}\left(\Delta c\mathrm{{}_{T,m}}-X\Delta\phi\mathrm{{}_{m}}\right)$ (45)
where $\Delta c\mathrm{{}_{T,m}}$ is the difference between $c\mathrm{{}_{T,m}}$ at the two outsides of the membrane, which we calculate based on Eqs. (34) and (35). For a 1:1 salt, we arrive at
@@
At high enough $|X|$ (relative to $c\mathrm{{}_{c}}$ and $c\mathrm{{}_{d}}$), we can derive from Eq. (46) that the difference in $c\mathrm{{}_{T,m}}$ across the membrane can be simplified to $\Delta c\mathrm{{}_{T,m}}=2\,\Phi_{i}^{2}\,\left(c_{\text{c}}^{2}-c_{\text{d}}^{2}\right){\big{/}}\,|X|$, a result that we can implement in Eq. (47) to arrive at
@@
which is the final equation that we need to solve the full co-current ED model. All fluxes and concentrations in Eq. (45) depend on position z in the module, or alternatively, depend on time, or time-on-stream, $t^{*}$.
@@
Eq. (48) shows that $\lambda\mathrm{{}_{th}}$ will be high, near unity, at high currents and a high membrane charge, and for low salt concentration differences, but it will go down for lower currents and lower membrane charge. It also goes down when $k^{*}_{\text{m}}$ goes up, for instance because membranes are thin. Indeed, for very thin membranes, ED does not work, because then there is a large flux of coions across the membranes and that limits the desalination process [43]. Eq. (48) illustrates that $\lambda\mathrm{{}_{th}}$ strongly depends on position in the cell, because current density I changes with z, as well as concentrations $c\mathrm{{}_{d}}$ and $c\mathrm{{}_{c}}$.
@@
Calculations with this model show that while water in the d-channel is desalinated, the Donnan potential strongly increases on the diluate side (by a value which exceeds the decrease on the concentrate side), the current goes down, and current efficiency decreases, finally to reach zero, see Fig. 5. In summary, the co-current ED model is a powerful and insightful model to discuss and quantify several well-known observations in ED, such as the decrease of the current density and current efficiency with time, see Fig. 5.
@@
Figure 5: Calculations of desalination in electrodialysis with co-current flow in diluate and concentrate channels for $V\mathrm{{}_{cp}}\!=\!180$ mV, $c\mathrm{{}_{f}}\!=\!0.5$ M, $\text{{wr}}\!=\!0.5$, $X\!=\pm\!4.0$ M, $k_{\text{m}}^{*}\!=\!1$ $\mu$m/s, $\Phi_{i}\!=\!1$, and $k\mathrm{{}_{ch}}\!=\!5$ $\mu$m/s. As function of time-on-stream, t*, salt concentrations in the two channels change, first rapidly, then slowly, while current I goes down as well as current efficiency, $\lambda\mathrm{{}_{th}}$. For a high t*, there is still a flow of current across the channels and membranes, but the net salt transport is zero, thus $\lambda\mathrm{{}_{th}}\rightarrow 0$.
Explanation of coion leakage in ED. Let us describe the functioning of the ED cell pair in an alternative manner. With the current directed to the right, cations will flow rightward, and anions will flow leftward. In the bulk of the channels, for a symmetric system, this current is equally carried by cations, moving right, and anions that move left. Cations and anions can have velocities in this left-right direction that tend to zero (near a blocking membrane), but they do not flow ‘back’ at any point in the module. With AEMs and CEMs alternatingly placed in the stack, and with cations moving right, anions left, then the following will happen. One type of channel has a CEM on its right side. That membrane freely allows the cations to pass to the next channel (to the right). And that same channel has an AEM on its left side, and anions can pass that membrane freely and move left. But now, in these adjacent channels, things are different. The anions want to continue moving left, but now they encounter a CEM which blocks them. And cations that want to move right, encounter an AEM that blocks them. So in this second set of channels, the ions accumulate, and thus the salt concentration increases. These channels will be the concentrate channels. Ions that arrive in these channels, cannot escape.
What was just described, is the ideal functioning of an ED stack, and it is generally like this early on in a cell pair, where current densities are still high and concentration differences between the two types of channels, $c\mathrm{{}_{d}}$ and $c\mathrm{{}_{c}}$, still low.1 But further on in the stack, with $c\mathrm{{}_{c}}\gg c\mathrm{{}_{d}}$, ions start to leak from the concentrate channel. They leak in the direction that drives them onward, not back to the diluate channel from which they came, but on to the next diluate channel. The membrane that blocked them from leaving the concentrate channel in the ideal case (e.g., the AEM for a cation), becomes leaky. Why is that? Chemically the membrane is of course the same as before. But this increased coion leakage from the concentrate channels is because the increasing concentration difference across this membrane leads to more and more diffusion of coions. Indeed, diffusion is the main membrane transport mechanism for co-ions in ED. Thus further on in a cell pair, the driving force for diffusion increases, and thus ions that were kept in the concentrate channel quite well at first, now increasingly leak out; so cations that entered a concentrate channel on the left side, now leak out on the right side.
@@ 5.1 Introduction
For transport across the membrane, be it for membranes as thin as 100 nm as in RO, or membranes 100 to 1000 times thicker for ED, we use continuum theory for flow of water and ions applied to transport in the pores in RO and ED membranes. The space charge (SC) model includes profiles in concentration and potential across the cross-section of a pore in the membrane, and also includes the effect of pressure on both membrane interfaces [36, 45, 50, 51, 46, 47, 48, 49]. But we use a simplification of the SC-model where concentrations are averaged over the cross-section of pore, which is the Uniform Potential (UP) model. The UP-model is an extension of the Teorell-Meyer-Sievers (TMS) model that describes ion transport across membranes, but TMS theory does not include water flow [52]. The UP-model can now be extended to incorporate many more effects than we could in the SC model. Besides water-membrane and water-ion friction, we can now also include ion-membrane friction and ion-ion friction, the latter describing how ions that overtake slower ones drag these along, and vice-versa slow ions will retard faster ones. This is the solution-friction (SF) model, and it describes both the flow of water and of solutes [53, 54]. In an extended version, also ion volume effects can be included, and this is called the two-fluid model (TFM) [55, 56, 14].
In SF theory, the velocity of solutes (such as ions) is determined by a balance between driving forces and frictions. Driving force acting on a solute are mathematically formulated as the negative of a gradient of an ion’s chemical potential. Frictions experienced by a solute are with the water flowing through the pores, and with the membrane matrix, which acts as an immobile background structure. For the water a similar force balance is set up but driving forces acting on the water are hydrostatic and osmotic pressures gradients [42, 40]. Thus ions and other solutes are described in a very different manner from how the water that flows past the ions through the pores is described. This is one of the key points of SF theory [57, 58, 59]. This difference in treatment is required because water is a continuum fluid and fills up all available space in the porous medium (such as a membrane), and thus forms continuous liquid pathways. Ions and other solutes are discrete, dispersed entities, and they travel across these pathways of water. These different modes of transport for water and for solutes are essential elements of SF theory.